# TermSight: Making Service Contracts Approachable

### v1

@@ 3. Formative Study
To better understand the challenges readers face when trying to make sense of the Term of Service and their information needs, we conducted a formative study to answer the following research questions:
@@ 3.1. Methodology
3.1.1. Study Procedure:
The entire study took 40 minutes and participants were compensated 10\$. The study was approved by the Institutional Review Board (IRB) office at our organization. Participants initially completed a demographics form and reported their past experiences interacting with ToS. Following the form, participants were asked to imagine as if they were a first-time user of the assigned service registering for the service and were provided 15 minutes to go through the ToS. Participants were free to navigate to different policies within the ToS. While reading, participants were instructed to speak aloud any challenges they faced or thoughts about the ToS. After reading the ToS, we conducted semi-structured interviews with participants about their experiences and challenges with reading ToS. Lastly, we asked the participants to envision how computational supports could aid their reading. All interview sessions were transcribed. The semi-structured interview questions are listed in Appendix E.
@@ 3.1.2. Analysis
We conducted a reflexive thematic analysis on the interview transcripts to identify common challenges and themes. We followed the six phases of reflexive thematic analysis suggested by Braun and Clarke  (Braun and Clarke, 2006). The lead researcher thoroughly explored the data, noted interesting features, systematically coded the data, and iteratively compared the codes to generate the initial themes. Then, the lead researcher discussed with other members of the research team to make sure the themes fit the data and further developed the themes.
@@ 3.1.3. Materials
Participants were randomly assigned one of the 10 ToS in the study. We selected ToS from 10 services spanning social media, e-commerce, video streaming, and internet services (Social Media: Reddit, Facebook, Instagram, Twitter; E-commerce: Amazon, eBay; Video: Youtube, Netflix; General: Google, Yahoo). These 10 services were selected from the top 20 visited sites on semrush.com (2024/05) where we selected the top 4 social media services, top 2 E-commerce services, top 2 video platforms, and top 2 internet services. We over-sampled social media services because they are the most represented service category in the top 10 visited services. In the study, participants were first directed to the main ToS and were allowed to navigate across any other policies.
@@ 3.1.4. Participants
We recruited 20 participants through Prolific. Each participant was randomly assigned one of the 10 selected ToS. All participants were over 18, fluent in English, and located in the US. 14 participants self-identified as female and 6 as male. 3 participants were between the ages of 18-25, 9 were between 26-35, 4 were between 36-45, 3 were between 46-55, and 1 was 67. 3 participants had never read or skimmed a ToS before and 17 participants had read or skimmed at least 1 ToS before. We did not require participants to have prior experience in reading the ToS. All the participants either had not read or did not remember reading the ToS they were assigned.
@@ 3.2. RQ1: What information do readers want from a ToS?
3.2.1. Participants have diverse information needs based on usage of the service and personal value
All 20 participants described the need to know what control and rights the service had over the user, what the service was allowed to do, and how the user could be negatively impacted. Additionally, 12 participants highlighted the need to know their (as opposed to the service’s) rights and control. Furthermore, 4 participants expressed wanting to know what they were allowed or not allowed to do on the platform. Participants also noted more specific and diverse information that they cared about such as data collection and usage (11), intellectual property right over user content (8), purchasing and returns (7), account deletion by the service (5), arbitration and liabilities (3), content moderation (2), and exposure to offensive content or misinformation (1).
10 participants explicitly mentioned how their information needs and priorities depended on usage and personal values. For example, P1 described that “I don’t think any information is irrelevant. I just think some parts are more of a priority of knowing. If I happen to become a person who wants to create graphic designs, I would definitely utilize the Copyright portion of the ToS to ensure that I’m being compliant with [the platform’s] expectations.” On the other hand, P3, with a background in software, raised privacy concerns when using services for professional communication: “For Microsoft Teams and Gmail, I was using them for a collaborative project with a small company. I wanted to ensure that whatever data or information was shared would remain proprietary to us.”
@@ 3.3. RQ2: What are the challenges readers encounter?
Our findings revealed that participants encountered challenges at the contract (§3.3.1), document (§3.3.2), and phrase (§3.3.4) level. In addition, participants highlighted how existing navigational and reading affordances were often ineffective, misleading, and failed to support meaningful comprehension (§3.3.3).
@@
11 participants mentioned that it’s unclear to them what sub-policies are included when they are navigating the ToS because sub-policies are often hyperlinked throughout the documents. There often lack a centralized list of what policies are included as part of ToS. 7 participants described the challenge of navigating across these nested policies as getting ‘lost in a whole tree of information’ (P18).
All 20 participants described relying on policy names to help decide whether to explore a policy. Yet, 17 participants described that the names of the policies alone did not provide a clear mental model of what was in the sub-policies and whether there was important information they should know about. For example, P14 mentioned that: “When it’s hyperlinked like ‘our rules and policies’ and ‘privacy policy’, it’s difficult to determine if I need to know something from the link.” 8 participants explicitly expressed their reluctance to click into the linked polices due to the lack of awareness of relevant information and the challenge of finding relevant information from a hyperlinked document.
@@
17 out of 20 participants described their reading behavior as skimming the ToS to see if any information caught their eye. 6 participants noted that they were not sure what they should look for in the ToS and might have unintentionally missed important information. P12 further commented how it’s not always obvious if the information was relevant at first sight: “A lot of it seemed irrelevant until I started reading in a little bit more.”
18 participants highlighted the overwhelming length of the document, especially when considering the nested policies. These participants expressed concerns about missing important terms buried in the text. For example, P7 pointed out that “When all the important information gets hidden behind walls of text, it doesn’t really incentivize anyone to read it, and it gets hidden.” Reflecting P7’s comment, 13 participants described the text as visually dense and difficult to parse when skimming. These participants further noted how important information can be positioned at the very bottom of an extended policy that was hard to discover. For example, P10 noticed how information about returns and cancellations of an E-commerce platform was positioned under the 13th section of the ToS with the title ‘Additional Terms’.
Additionally, 3 participants described how the use of friendly language actually obfuscated the ToS and made skimming more challenging. For example, P15 noticed language such as ‘We only use data to make [the service] a better place’ when reading a privacy policy. Yet, she later found out how user information can be used for targeted ads. As a result, P15 described that it’s easy for lay people to misinterpret the language and miss important terms: “If someone’s just skimming the terms of service and is not skeptical at all. Then, you could read over that and think that means they’re not using your data for targeted ads or stuff.”
@@ 3.3.3. Document Level: Participants found existing affordances to be ineffective and misleading
Some of the ToS participants read contained navigational and reading affordances like an interactive table of contents with section headers and overview summaries. However, 12 participants noted that, similar to policy names, section headings within policies were vague and not descriptive of the actual content. As a result, it was unclear what information was contained in a section by only looking at the section header. For example, when skimming the ToS, P10 skipped the ‘Content’ section but later realized that the section was about intellectual property rights over user-generated content: “It just says ‘content’ which is pretty vague when they’re talking about getting exclusive intellectual rights. I feel like it’s misleading. So people would skip right past it. I did the first time.”
In addition to ineffective navigational affordances such as vague policy names (17/20) and section headers (12/20), 3 participants accessed a sub-policy with a summary at the top. All three participants (P4, P14, P15) described using the summaries to gain an overview at the start. Yet, after reading the policy, participants noticed that the summary was missing key details: “The summary is kind of misleading. They gave you a more palatable version of [ToS]. If you scroll down, it will say we can terminate your account for no reason. They don’t say that in the summary” (P4). They also expressed concern that readers might only glance at the summary and mistakenly think it covers the entire policy.
@@ 3.3.4. Phrase Level: Participants struggled to contextualize unfamiliar or ambiguous terminology
When reading individual sections in a policy, 12 participants pointed out unfamiliar legal terminology (e.g., Arbitration, Indemnity, or class lawsuit). These participants described how the terminology hindered their understanding and motivation to read the section. For example, P8 described skipping sections that were full of unfamiliar terms. Additionally, participants noted how challenging it was to interpret the meaning of a word or phrase in the context of their situation: “I can recognize these are legal terms, but I don’t necessarily know what that means for me” (P10). P6 also described how it’s hard to envision scenarios in which things go wrong beforehand from the ToS.
8 participants noticed ambiguous or vague terminologies that made it challenging for them to understand the implications of signing. For example, P15 mentioned that “I really think the big difficulty is the vagueness of the language and the constancy of exceptions that are vague which gives them a lot more leeway.” More specifically, 5 participants noticed vague terms around the service’s information collection and sharing practices such as: ‘we share you data with 3rd parties’, ‘track information about your interaction’, and ‘retain certain information about you’. 4 participants noted vague terms around how the service can use user-generated content. For example, P3 and P14 pointed out that it’s unclear what “use, distribute, modify, and copy user content” implicates and what are examples of what the service can do with user-generated content.
3.4. Design Guidelines:
@@
DG3: Phrase level: Help readers read and interpret original text with technical language, jargon, and vague phrases.
@@
4. The TermSight System:
We reify these design guidelines in TermSight, an augmented reading interface that supports user sense-making of ToS with features at each level of granularity of a ToS:
@@ 4. The TermSight System:
Summary Snippets: 1-sentence plain language summaries of information chunks at the document-level. Snippets are rendered with color and saturation that represent power and relevance to the user, similar to the Power Meter.
@@
We adopted an iterative design approach in developing TermSight. 8 participants evaluated an early prototype of TermSight, and their feedback informed the final design. Additional details about the iterative design study, results, and design changes can be found in Appendix  A.
4.1. Design Abstractions:
The design of TermSight is grounded in two design abstractions: 1) Information Snippets, and 2) Power and Relevance classification. These abstractions serve as foundational building blocks for TermSight’s features that support contract and document-level sensemaking as illustrated in Figure  6.
@@ 4.1.1. Information Snippets
Our formative study reveals that the information readers cared about in ToS is often buried within extended and visually dense text. Consequently, prior works that support document understanding through document-level summaries (Tomuro et al., 2016; Keymanesh et al., 2020) or section-level summaries (Manor and Li, 2019; August et al., 2023) can fall short in helping readers surface pieces of information buried in large sections. To address this limitation, we use Information Snippets as a basic building block for TermSight. An information snippet is a continuous span of original text that shares the same topic and can be summarized in a single sentence of up to 12 words (detailed in Section 5.2). These information snippets form the basis of the Summary Snippets (§4.3.1). Information Snippets are non-overlapping and together reconstruct the original document’s content (Examples in Figure  2). Figure  6 illustrates the process of how TermSight divides a paragraph into five Information Snippets, each paired with its Summary Snippet. The idea of designing interactions and sense-making around Information Snippets rather than entire sections was inspired by prior works that modularize and ‘objectify’ tools (Bier et al., 1993; Bederson et al., 1996), attributes (Xia et al., 2016), and AI agent’s memory (Huang et al., 2023) to enable interaction flexibility, interaction specificity, and direct manipulation.
Figure 2. Design conceptualization of Power and Relevance: Power refers to the degree to which a snippet grants control to the service provider or the user. Relevance refers to whether or not the snippet is relevant to the user’s intended usage of the service and personal value. Examples of Information Snippets, TermSight classified under each category of power, are shown.
@@ 4.1.2. Power and Relevance Classification
We define Power and Relevance as two qualities of an information—and thus a summary—snippet. Power is the degree to which a snippet’s text grants control to the Service Provider or the User (Categories: Service, Neutral, or User). Relevance is whether the snippet’s text is directly relevant to the user’s intended usage of the service or their values (Categories: High, Low). For example, a snippet related to selling would have low relevance for a user more focused on buying. As shown in Figure  2, we visually encode Power and Relevance in TermSight with a combination of hue (Red, Yellow, Green) and saturation (High, Low).
@@
Figure 3. Power Meter: a visualization of the distribution of the power and relevance of Information Snippets within a document (a). Users can gain an overview of the policies within the ToS (b). When hovering over the Power Meter, the list of Summary Snippets of the corresponding policy is shown (c). Clicking the blue backquote would allow users to view the referenced Information Snippet (d).
4.2. Power Meter: Policy level visualization of power and relevance
While prior works have focused on a single policy such as the Privacy Policy  (Nokhbeh Zaeem et al., 2022; Tesfay et al., 2018) or the main ToS page (Lippi et al., 2019; Guarino et al., 2021), Terms of Services include many policies (e.g., content policy, community guidelines, return policy, and more). To help users form mental models of each sub-policy, TermSight provides information scent and preview to help users gain an overview of the policies within ToS to decide which policies to read (DG1).
4.2.1. Providing information scent of a policy
In TermSight, the policies that are part of the ToS agreement are placed in the top navigation panel where users can navigate to different policies. Each policy in the top navigation panel or hyperlinked within the document is accompanied by a Power Meter: a horizontal bar visualization representing the distribution of Information Snippets in a document, with colors denoting power and relevance. Shown in Figure 3, Power Meter provides information scent (Pirolli and Card, 1999) of the power distribution of Information Snippets within the policy and the fraction of Information Snippets that are potentially relevant to the user.
4.2.2. Providing a preview of the document
Hovering over a Power Meter reveals a preview popup containing a list of Summary Snippets, allowing users to skim and gain an overview of a policy’s Information Snippets without having to be redirected to a new page (Figure 3c). Each Summary Snippet in the popup includes a clickable quote icon that allows users to view the referenced Information Snippet (DG4).
4.3. Summary Snippets: Plain language summaries of discrete information snippets

### v2

@@ 3. Formative Study
We first establish the challenges general-audience readers face when reading a legal contract. We focus on Terms of Service (ToS) because ToS are legally binding contracts that are signed routinely in digital interactions. While ToS are rarely read completely (Obar and Oeldorf-Hirsch, 2020; Good et al., 2005; Ahmad et al., 2020), past legal work suggests that these contracts represent many of the broader issues faced in legal contracts today (Obar and Oeldorf-Hirsch, 2020; Kar and Radin, 2019) and therefore present an ideal setting to explore possible reading features. Our formative study was guided by the following research questions:
@@ 3.1. Methodology
3.1.1. Study Procedure
The entire study took 40 minutes, and participants were compensated 10\$. The study was approved by the Institutional Review Board (IRB) office at our organization. Participants first reported their past experiences interacting with ToS. Next, participants were asked to imagine as if they were a first-time user registering for a randomly assigned service and were provided 15 minutes to go through the ToS while speaking aloud any challenges. Participants were free to navigate to different policies within the ToS. After reading the ToS, we conducted semi-structured interviews about participants’ experiences and challenges with ToS reading, along with opportunities for computational support. The interview questions are listed in Appendix D.
@@ 3.1.2. Analysis
We conducted a reflexive thematic analysis on the interview transcripts to identify common challenges and themes. We followed the six phases of reflexive thematic analysis suggested by Braun and Clarke  (Braun and Clarke, 2006). The lead researcher thoroughly explored the data, noted interesting features, systematically coded the data, and iteratively compared the codes to generate the initial themes. Through weekly meetings over the course of two months, the lead researcher discussed with other members of the research team to make sure the themes fit the data and further developed the themes. This included iterating on the descriptions of the themes and merging related themes.
@@ 3.1.3. Materials
Participants were randomly assigned one of the 10 services’ ToS [^2] Social Media: Reddit, Facebook, Instagram, Twitter; E-commerce: Amazon, eBay; Video: Youtube, Netflix; General: Google, Yahoo. The 10 services were selected from the top 20 visited sites on semrush.com (2024/05), where we sampled the top 4 social media services, top 2 E-commerce services, top 2 video platforms, and top 2 internet services. We oversampled social media services because they were the most represented service category in the top 10 visited services.
@@ 3.1.4. Participants
We recruited 20 participants through Prolific (Male: 6, Female: 14). All participants were over 18, fluent in English, and located in the US. Participants’ age ranged from 20 to 67 ($\mu=35.75,\sigma=12.24$). 3 participants had never read or skimmed a ToS, and 17 participants had read or skimmed at least 1 ToS before. We did not require participants to have prior experience in reading the ToS. All the participants had not read or did not remember reading the ToS they were assigned.
@@ 3.2. RQ1: What information do readers want from a ToS?
3.2.1. Participants have diverse, value-dependent information needs when reading ToS
All 20 participants described the need to know what control and rights the service had over the user, what the service was allowed to do, and how the user could be negatively impacted. While all participants shared a desire to know their general rights, different participants also noted more specific information needs, such as data collection and usage (11), intellectual property rights (8), purchasing and returns (7), account deletion by the service (5), arbitration and liabilities (3), content moderation (2), and exposure to offensive content or misinformation (1). 10 participants explicitly mentioned how their information needs from the ToS and priorities depended on service usage and personal values. For example, P1 described that “I don’t think any information is irrelevant. I just think some parts are more of a priority of knowing. If I happen to become a person who wants to create graphic designs, I would definitely utilize the Copyright portion of the ToS to ensure that I’m being compliant with [the platform’s] expectations.” On the other hand, P3, with a background in software, raised privacy concerns when using services for professional communication: “For Microsoft Teams and Gmail, I was using them for a collaborative project with a small company. I wanted to ensure that whatever data or information was shared would remain proprietary to us.”
@@ 3.3. RQ2: What are the challenges readers encounter?
Our findings revealed that participants encountered challenges at the contract (§3.3.1), document (§3.3.2), and phrase (§3.3.4) level. These challenges reflected the complexity and ambiguity of legal language. In addition, participants highlighted how existing navigational and reading affordances were often misleading and failed to support meaningful sensemaking of the actual ToS text (§3.3.3).
@@
Eleven participants mentioned that it’s unclear to them what sub-policies are included when they are navigating the ToS because sub-policies are often hyperlinked throughout the documents. There often lacks a centralized list of what policies are included as part of ToS. Seven participants described the challenge of navigating across these nested policies as getting ‘lost in a whole tree of information’ (P18). All 20 participants described relying on policy names to help decide whether to explore a policy. Yet, 17 participants noted that the names of the policies alone did not provide a clear mental model of what was in the sub-policies and whether there was important information they should know about. For example, P14 mentioned that: “When it’s hyperlinked like ‘our rules and policies’ and ‘privacy policy’, it’s difficult to determine if I need to know something from the link.” 8 participants explicitly expressed their reluctance to click into the linked polices due to the lack of awareness of relevant information and the challenge of finding relevant information from a hyperlinked document.
@@
Most (17/20) participants primarily skimmed the ToS. Six participants noted that they were not sure what they should look for in the ToS and might have unintentionally missed important information. P12 further commented how it’s not always obvious if the information was relevant at first sight: “A lot of it seemed irrelevant until I started reading in a little bit more.” Often, this inability to find information was connected to the overwhelming length of the document (18/20). Interestingly, many (13/20) participants also noted how text positioning seemed intentionally obfuscating: important information was positioned at the bottom of an extended policy. For example, P10 noticed how information about returns and cancellations of an E-commerce platform was positioned under the 13th section with the title ‘Additional Terms’. Similarly, three participants described how the use of friendly language tended to obfuscate information and made skimming more challenging. For example, P15 noticed language such as ‘We only use data to make [the service] a better place’ when reading a privacy policy. Yet, she later found out how user information can be used for targeted ads.
@@ 3.3.3. Document Level: Participants found existing affordances to be ineffective and misleading
Some of the ToS participants read contained navigational and reading affordances, such as interactive table of contents with section headers and overview summaries. However, 12 participants noted that, similar to policy names, section headings within policies were vague and not descriptive of the actual content. For example, when skimming the ToS, P10 skipped the ‘Content’ section but later realized that the section was about intellectual property rights over user-generated content: “It just says ‘content’ which is pretty vague when they’re talking about getting exclusive intellectual rights. I feel like it’s misleading. So people would skip right past it. I did the first time.” In addition to ineffective navigational affordances such as vague policy names (17/20) and section headers (12/20), three participants accessed a sub-policy with a summary at the top. All three participants (P4, P14, P15) described using the summaries to gain an overview at the start. Yet, after reading the policy, participants noticed that the summary was missing key details: “The summary is kind of misleading. They gave you a more palatable version of [ToS]. If you scroll down, it will say we can terminate your account for no reason. They don’t say that in the summary” (P4).
@@ 3.3.4. Phrase Level: Participants struggled to contextualize unfamiliar or ambiguous terminology
When reading individual sections in a policy, 12 participants pointed out unfamiliar legal terminology (e.g., Arbitration, Indemnity, or class lawsuit). These participants noted the difficulty to interpret the meaning of legal jargon in the context of their situation: “I can recognize these are legal terms, but I don’t necessarily know what that means for me” (P10). In a similar vein, eight participants noticed ambiguous terminology that made it challenging for them to understand the implications of signing (e.g., ‘we share your data with 3rd parties’, and ‘retain certain information about you’). P15 mentioned that “I really think the big difficulty is the vagueness of the language and the constancy of exceptions that are vague, which gives them a lot more leeway”.
3.4. Design Guidelines
@@
DG3: Phrase level: Help readers read and interpret original text with technical language, jargon, and ambiguous phrases.
@@
4. The TermSight System
We reify these design guidelines in TermSight (Figure  1), an augmented reading interface for exploring opportunities to make ToS contracts more approachable with features at every granularity:
@@ 4. The TermSight System
Summary Snippets: One-sentence plain language summaries of information chunks at the document-level. Snippets are rendered with color and saturation that represent power and relevance to the user, similar to the Power Meter.
@@
We adopted an iterative design approach in developing TermSight. Eight participants evaluated an early prototype of TermSight, and their feedback informed the final design. Additional details about the iterative design study, results, and design changes can be found in Appendix  A.
4.1. Design Abstractions
The design of TermSight is grounded in two design abstractions: 1) Information Snippets, and 2) Power and Relevance classification.
@@ 4.1.1. Information Snippets
Our formative study revealed that the information readers cared about in ToS can be buried within visually dense text. To enable sense-making of finer-grained pieces of information as opposed to an entire section, we use Information Snippets as the unit of interaction in TermSight. An Information Snippet is a continuous span of original text that shares the same topic and can be summarized in a single sentence of up to 12 words (detailed in Section 5.2). These Information Snippets form the basis of the Summary Snippets (§4.3.1). Information Snippets are non-overlapping and together reconstruct the original document’s content. Figure  5 illustrates the process of how TermSight divides a paragraph into five Information Snippets, each paired with its Summary Snippet. The idea of designing interactions and sense-making around finer-grained Information Snippets was inspired by prior works that modularize and ‘objectify’ tools (Bier et al., 1993; Bederson et al., 1996), attributes (Xia et al., 2016), and AI agent’s memory (Huang et al., 2023) to enable interaction specificity and direct manipulation.
@@ 4.1.2. Power and Relevance Classification
Our formative study revealed participants’ information needs regarding the rights and control held by the service and the user. These needs may also vary depending on the intended service usage. To provide information scent  (Pirolli and Card, 1999) for snippets of information that a reader may care about, we reify participants’ information needs by defining Power and Relevance as two qualities of an information—and thus a summary—snippet. Power is the degree to which a snippet’s text grants control to the Service Provider or the User (Categories: Service, Neutral, or User). Relevance is whether the snippet’s text is directly relevant to the user’s intended usage of the service or their values (Categories: High, Low). For example, a snippet related to selling would have low relevance for a user more focused on buying. As shown in Figure  2, we visually encode Power and Relevance in TermSight with a combination of hue (Red, Yellow, Green) and saturation (High, Low). We constrained relevance to High and Low because our iterative design study revealed that more color–saturation pairs (i.e., 3 hues for Power x 3 saturations for Relevance) were difficult to interpret (Appendix  A).
Figure 2. Design conceptualization of Power and Relevance. Power refers to the degree to which a snippet grants control to the service provider or the user. Relevance refers to whether or not the snippet is relevant to the user’s persona.
The figure shows how TermSight defines and visualizes power and relevance using color. Power indicates the degree to which a snippet grants control to the service provider or the user (red = favor provider, yellow = neutral, green = favor user). Relevance indicates whether a snippet is pertinent to the user’s persona, which is based on intended service usage and personal values (low saturation = low relevance, high saturation = high relevance). The figure also provides examples. A provider-favoring clause: ‘This license includes the right for us to use your content to train machine learning models…’; A neutral clause: ‘Avoid posting illegal content’; and a user-favoring clause: ‘Our licensees cannot display public content deleted by the users…’.
4.2. Power Meter: Visualization of the Distribution of Power and Relevance
In TermSight, the policies that are part of the ToS are placed in the top navigation panel. To help users form mental models of each policy and decide which policies to read (DG1), each policy in the top navigation panel or hyperlinked within the document is accompanied by a Power Meter: a horizontal bar visualization representing the distribution of Information Snippets in a document, with colors denoting power and relevance (Figure 3). Our iterative design study further revealed the need to contextualize Power Meter with concrete previews (Appendix  A). Consequently, hovering over a Power Meter reveals a policy preview containing a list of Summary Snippets (Figure 3b). Each Summary Snippet in the popup includes a clickable icon (Figure 3c) that allows users to view the referenced Information Snippet (DG4).
Figure 3. Power Meter visualizes the distribution of power and relevance of Information Snippets within a policy (a). On hover, a preview of Summary Snippets is shown (b) with options to view the referenced Information Snippets (c).
@@ 4.2. Power Meter: Visualization of the Distribution of Power and Relevance
The figure visualizes the Power Meter. (a) The Power Meter is a horizontal bar that visualizes the distribution of power and relevance of Information Snippets within a policy. (b) When a user hovers over the Power Meter, the system displays the list of Summary Snippets for that policy. (c) Each Summary Snippet is paired with a blue backquote icon that links to the referenced Information Snippet.
4.3. Summary Snippets: Plain language Summaries of Information Snippets